OpenStudy (anonymous):

For this question, you will need a parent/guardian or a friend. Have this individual grab a handful of coins making sure there are only two types of coins in the group (i.e., nickels and dimes, quarters and pennies, pennies and dimes, etc). Your parent/guardian or friend should tell you the type of coins they’ve chosen, how many coins they have and the dollar amount of the group. From this information, you will set up two sets of equations and determine how many of each coin they have in their hand. Please send your instructor the name of the individual who helped you with this question, your your two equations and the work you did to solve the system.

OpenStudy (anonymous):

@Rexy705 can u help?

OpenStudy (anonymous):

you want me to make something up?

OpenStudy (anonymous):

i have 9 coins only dimes and quarters the total amount of money is $1.65 how many of each do i have?

OpenStudy (anonymous):

idk i need help @satellite73

OpenStudy (anonymous):

lol @question it implies that those whose parent/guardian is no longer alive can't do the question. How terribly restricting..

OpenStudy (anonymous):

ok well there is the first part i came up with a fictitious amount of dimes and quarters and a total amount of money now comes the second part, you have to figure out how many of each i have

OpenStudy (anonymous):

lol @genius12 i guess i am a parent for a day

OpenStudy (anonymous):

ok how?

OpenStudy (anonymous):

lets put \(x\) as the number of dimes, \(y\) as the total number of quarters, and solve for \(x\) and \(y\)

OpenStudy (anonymous):

i told you i had 9 coins all together how can we write that as an equation with \(x\) and \(y\) ?

OpenStudy (anonymous):

x+y=9?

OpenStudy (anonymous):

got it! first equation is done

OpenStudy (anonymous):

now for the second one each dime is worth 10 cents, so if i have \(x\) dimes, how much money do i have in dimes (hint, if i had 13 dimes i would have 130 cents)

OpenStudy (anonymous):

if it is not clear, let me know and i will tell you

OpenStudy (anonymous):

90?

OpenStudy (anonymous):

if i had 9 dimes i would have 90 cents, yes, but i don't have 9 dimes, i have \(x\) dimes your answer has to have an \(x\) in it

OpenStudy (anonymous):

what did you do to 9 to get 90?

OpenStudy (anonymous):

like i said, if it is not clear, let me know and i will tell you

OpenStudy (anonymous):

did i lose you?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is clear

OpenStudy (anonymous):

ok so is it more or less clear that the amount of money i have in times is \(10x\) ?

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

if i have 3 dimes i have 30 cents if i have 6 dimes i have 60 cents if i have 9 dimes i have 90 cents if i have \(x\) dimes i have \(10x\) cents

OpenStudy (anonymous):

ohhhhh okay

OpenStudy (anonymous):

similarly if i have 2 quarters i have 50 cents if i have 4 quarters i have 100 cents if i have 7 quarters i have 175 cents

OpenStudy (anonymous):

how much if i have \(y\) quarters?

OpenStudy (anonymous):

10y?

OpenStudy (anonymous):

not quite \(10\) is for the dime how many pennies in a quarter?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

25y?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so now the total amount of money: \(10x\) for the dimes, \(25y\) for the quarters, so the total is \[10x+25y\] which we know is \(165\) so we can say \[10x+25y=165\]

OpenStudy (anonymous):

now we have two equations \[x+y=9\\10x+25y=165\] and we can solve this for \(x\)

OpenStudy (anonymous):

can you help me solve?

OpenStudy (anonymous):

any idea how to do it? ("no" is a find answer, i will show you if you do not)

OpenStudy (anonymous):

ok i take that as a no

OpenStudy (anonymous):

the first equation \(x+y=9\) is the same as saying \(y=9-x\)

OpenStudy (anonymous):

then we can replace the \(y\) in the second equation \(10x+25y=165\) by \(\color{red}{9-x}\) and solve for \(x\) \[10x+25\color{red}y=165\\ 10x+25(\color{red}{9-x})=165\]

OpenStudy (anonymous):

multiply out on the left to get \[10x+225-25x=165\]

OpenStudy (anonymous):

collect like terms on the left to get \[-15x+225=165\]

OpenStudy (anonymous):

subtract \(225\) from both sides and get \[-15x=165-225=-60\] and finally divide both sides by \(-15\) and get \[x=(-60)\div( -15)=4\]

OpenStudy (anonymous):

so i have \(4\) dimes now the total was \(9\) so how many quarters do i have?

OpenStudy (anonymous):

ummm 4?

OpenStudy (anonymous):

total is 9 4 dimes how many quarters?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i don't believe you

OpenStudy (anonymous):

i"m confused

OpenStudy (anonymous):

9 coins, 4 dimes, how many quarters

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because \(9=4+5\)

OpenStudy (anonymous):

now lets check the answer 5 quarters and 4 dimes is 9 coins the 5 quarters are worth $1.25 and the 4 dimes are worth $.40 for a total of $1.25+.40=1.65

OpenStudy (anonymous):

oh thanks that helped a lot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

where is says Please send your instructor the name of the individual who helped you with this question you can put "satellite73"

OpenStudy (anonymous):

lol sure